OpenStudy (anonymous):

prove the identity: (1+cosx)(cox+sinx-1)=(sinx)(cosx-sinx+1)

OpenStudy (amoodarya):

(1+cosx)(cox+sinx-1) = 1(cox+sinx-1) + cos x (cox+sinx-1) =cos x+ sin x -1 + cos^2 x +cos x sin x -cos x cos x cancel = sinx -1 +cos ^2x +cos x sin x = sin x -(1-cos^2x) +cos x sin x = sin x - sin ^2x + cos x sin x now factor sin x = sin x (1-sin x + cos x ) it is rs